OpenStudy (anonymous):

The potential difference across a resistor is 12V. The current in the resistor is 2.0A. 4.0 C passes through the resistor. What is the energy transferred and the time taken?

OpenStudy (anonymous):

The definition of ampere is "coulombs per second", and therefore the current will last 2 seconds. The power in resistor is P=V*I=12 V * 2 A = 24 watts. Therefore, the energy transferred is 24 W * 2 s = 48 joules.